Question 5
a.   Given the following cash flows, for the two independent projects A and B, calculate
i.   Payback Period   
ii.   Accounting rate of return   
iii.   Net Present Value
iv.   Profitability index
And recommend acceptance or rejection of projects considering individual techniques of capital budgeting. A rate of 10 % has been selected for the NPV analysis.

    Project A   Project B
Initial outlay   $50,000   $100,000
Cash inflows        
Year 1   $10,000   $ 25,000
Year 2   15,000   25,000
Year 3   20,000   25,000
Year 4   25,000   25,000
Year 5   30,000   25,000
  

b.   Explain the distinctive features of capital budgeting decisions.                                  
Answer #1

as per HomeworkLib policy please find below answer of 1st question (which has 4 sub part) for another q raise new request.

Cumulative cash flow
year Project A Project B Project A Project B
0 -50000 -100000 -50000 -100000
1 10000 25000 -40000 -75000
2 15000 25000 -25000 -50000
3 20000 25000 -5000 -25000
4 25000 25000 20000 0
5 30000 25000 50000 25000
i.   Payback Period   
Project A = =2+5000/25000 2.2 year
Project B = 4 year
ii.   Accounting rate of return   
Project A
Average profit = =50000/5 10000
Average investment = 50000= 50000
ARR = 10000/25000= 20%
Project B
Average profit = =25000/5 5000
Average investment = 100000= 100000
ARR = 5000/50000= 5%
iii.   Net Present Value
Present value
i ii iii iv=i*iii v=ii*iii
year Project A Project B PVIF @ 10% Project A Project B
0 -50000 -100000     1.0000            (50,000)             (100,000)
1 10000 25000     0.9091                9,091                 22,727
2 15000 25000     0.8264              12,397                 20,661
3 20000 25000     0.7513              15,026                 18,783
4 25000 25000     0.6830              17,075                 17,075
5 30000 25000     0.6209              18,628                 15,523
Total              22,217                 (5,230)
NPV A $ 22,216.88
NPV B $ (5,230.33)
iv.   Profitability index
Profitability index = present value of cash flow / Initial investment
PI of project A = =(50000+22217)/50000                  1.44
PI of project B = =(100000-5230)/100000                  0.95
Decision criterial =
Payback
Project A = Accepted It has lower payback period
Project B = Rejected It has higher paybak period
ARR
Project A = Accepted It has higher ARR
Project B = Rejected It has lower ARR
NPV
Project A = Accepted Since NPV is positive
Project B = Rejected Since NPv is negative
PI
Project A = Accepted PI is > 1
Project B = Rejected PI is < 1